# Memento-II: Learning by Stateful Reflective Memory

### v1

@@ 3.4 Two-Time-Scale Convergence with Online Memory Re-writing
    $\sum_{t=0}^{\infty}\eta_{t}=\infty,\quad\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty,\quad\sum_{t=0}^{\infty}\rho_{t}=\infty,\quad\sum_{t=0}^{\infty}\rho_{t}^{2}<\infty,\quad\lim_{t\to\infty}\frac{\rho_{t}}{\eta_{t}}=0$
@@
    $\mathbb{E}[\xi^{(Q)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[\xi^{(\mu)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[(\xi^{(Q)}_{t})^{2}|\mathcal{F}_{t}]\leq K,\quad\mathbb{E}[(\xi^{(\mu)}_{t})^{2}|\mathcal{F}_{t}]\leq K$
@@
    Memory attractor: The memory process $\{M_{t}\}$ has a compact attractor set $\mathcal{M}_{\infty}\subset\mathcal{M}$. This means memory size is bounded and the content distribution stabilises over time. In practice, this is achieved by: 1) Using fixed-size memory, 2) Having policies that converge to stationary distributions, 3) Using environments that produce stationary experience streams, or 4) Conservative memory update strategies.
@@
Under Assumptions 5 and 8, let $(Q_{t},\mu_{t},M_{t})$ be generated by the following sample-based recursions.
@@
The above result can be established using the ODE method for stochastic approximation [borkar2009stochastic]. The key insight is that from the perspective of the fast variables $(Q_{t},\mu_{t})$, the slow variable $M_{t}$ appears quasi-static due to the timescale separation $\rho_{t}/\eta_{t}\to 0$. This detailed proof establishes several key properties of our Read–Write Learning iteration. First, it demonstrates robustness to memory updates, showing that the algorithm remains stable even as the memory evolves, provided that memory updates occur on a slower timescale than value and policy updates. Second, it guarantees tracking behaviour, ensuring that the value function and policy rapidly adjust to reflect changes in memory, remaining close to the optimal configuration for the current memory state. Third, the analysis confirms convergence to local optima, proving that for each fixed memory configuration (M), the algorithm converges to the corresponding optimal retrieval policy ($\mu^{*}_{M}$). Finally, the timescale separation result provides practical implementation guidance, indicating that memory should be updated less frequently than the value function and policy. Collectively, these results provide strong theoretical guarantees for continual Read–Write Learning loop under our SRDP, demonstrating that it can adapt effectively to evolving memory while preserving convergence and stability.
@@ 3.5 The Role of Reflection and Memory towards Optimal Policy
A key question now is whether the Read–Write Learning agent can approach optimal performance in the underlying MDP as memory grows. We now establish reflection consistency: the property that increasing memory coverage drives the composite policy toward the true optimal policy.
@@
Algorithm 2 Memory-Based Policy Iteration via Parzen–KL Q-Learning
1:Input: Learning rate $\eta_{t}$, temperature $\alpha$, bandwidth $h$
2:Initialise: Episodic memory $M_{0}$ and Q-function $Q_{0}$
3:
4:(1) Read: Retrieve relevant experiences
5:Given current augmented state $x_{t}=(s_{t},M_{t})$, sample a case $c_{t}\sim\mu(\cdot\mid x_{t})$ from memory.
6:(2) Evaluate: Compute soft target using Parzen–KL smoothing
7:Observe reward $r_{t}$ and next augmented state $x_{t+1}=(s_{t+1},M_{t+1})$
8:Estimate Parzen prior $\mu_{0}(\cdot\mid x_{t+1})$ using (6)
9:Compute the soft target:
$y_{t}=r_{t}+\gamma\alpha\log\sum_{c^{\prime}\in M_{t+1}}\mu_{0}(c^{\prime}\mid x_{t+1})\exp\!\big(Q_{t}(x_{t+1},c^{\prime})/\alpha\big)$
10:(3) Write: Update memory and value function
11:Update Q-value:
$Q_{t+1}(x_{t},c_{t})=(1-\eta_{t})Q_{t}(x_{t},c_{t})+\eta_{t}y_{t}$
12:Store the new experience $(s_{t},c_{t},r_{t},s_{t+1})$ into memory:
$M_{t+1}=\mathsf{Write}(M_{t},s_{t},c_{t},r_{t})$
13:
14:Output: Updated memory $M_{t+1}$ and Q-function $Q_{t+1}$
For each memory $M$, define the coverage radius $r_{M}\triangleq\sup_{s\in\mathcal{S}}\min_{c\in M}d\big(s,s(c)\big)$. Let the (deterministic) retrieval rule $c_{M}(s)$ pick any case achieving the minimum. The LLM-induced composite policy on the original state space is
$\pi_{M}(a\mid s)\;\triangleq\;p_{\mathrm{LLM}}\big(a\mid s,c_{M}(s)\big).$
@@
Lemma 11 (Policy Approximation at State Level).
@@
$\delta_{M}\;\triangleq\;\sup_{s\in\mathcal{S}}\delta_{M}(s).$
@@
$\Delta_{M}\;\triangleq\;\sup_{s\in\mathcal{S}}\mathrm{TV}\!\big(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\big)\;\leq\;\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}.$
@@
4 Read–Write Learning Equivalency
The following theorem shows that iterative memory operations, consisting of the read and write processes, are equivalent to performing policy iteration, which includes policy evaluation and policy improvement. This demonstrates that memory interaction itself realises a learning mechanism, and by the previous theorem, the process converges to the optimal policy in the underlying MDP.
Theorem 15 (Equivalence of Memory and Policy Operations).
The following processes are equivalent:
  1. 1.
    Retrieval Policy Iteration in SRDP:
    • Memory Read: $c\sim\mu(\cdot|x)$
    • Memory Write: $M^{\prime}=\mathsf{Write}(M,s,a,r)$
    • LLM Action: $a\sim p_{\mathrm{LLM}}(\cdot|s,c)$
  2. 2.
    Standard Policy Iteration in Reflected MDP $\widetilde{\mathcal{M}}$:
    • Policy Evaluation: Compute $Q^{\mu}(x,c)$
    • Policy Improvement: Update $\mu$ to be greedy/softmax w.r.t. $Q^{\mu}$
Specifically:
  1. (a)
    The value functions are equal: $V^{\pi_{M}}(s)=\widetilde{V}^{\mu}(x)$
  2. (b)
    Optimal policies correspond: $\pi_{M}^{\star}$ optimal in M-MDP $\iff$ $\mu^{\star}$ optimal in $\widetilde{\mathcal{M}}$
  3. (c)
    Memory write implements policy evaluation data collection
  4. (d)
    Memory read implements policy improvement
Proof.
(a) Value Function Equality:
$V^{\pi_{M}}(s)$ $=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\middle|s_{0}=s,a_{t}\sim p_{\mathrm{LLM}}(\cdot|s_{t},c_{t}),c_{t}\sim\mu(\cdot|x_{t})\right]$
$=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\widetilde{\mathcal{R}}(x_{t},c_{t})\middle|x_{0}=(s,M_{0}),c_{t}\sim\mu(\cdot|x_{t})\right]$
$=\widetilde{V}^{\mu}(x)$
(b) Policy Optimality: Follows directly from (a) since the objectives are identical.
(c) Memory Write as Evaluation: Storing $(s,a,r,s^{\prime})$ provides the data $(x,c,\widetilde{\mathcal{R}}(x,c),x^{\prime})$ needed for temporal difference learning.
(d) Memory Read as Improvement: Selecting $c$ to maximise expected return is exactly policy improvement in $\widetilde{\mathcal{M}}$. ∎
Therefore, in practical implementations, memory operations can be used as a mechanism for continuously learning and improving the policy over time, as illustrated in Algorithm 2.
@@ 5 Related Work
Our work on Memory-Augmented Markov Decision Processes (LLM-M-MDPs) with Parzen-KL retrieval builds upon, yet also departs from, several established lines of research in reinforcement learning and large language models (LLMs). This section situates our contribution within the broader landscape of memory-based RL and clarifies the key distinctions from existing approaches.
The integration of memory mechanisms into reinforcement learning has long been explored to address the challenges of partial observability and long-term credit assignment. Early approaches often relied on recurrent neural networks (RNNs), in particular Long Short-Term Memory (LSTM) networks [hochreiter1997long], to maintain hidden states that capture historical information. While effective in certain tasks, these methods frequently struggle with long-term dependencies and lack explicit mechanisms for memory management [morad2023memory].
@@
Episodic control methods form another strand of memory-based RL. Model-Free Episodic Control (MFEC) [blundell2016model] and Neural Episodic Control (NEC) [pritzel2017neural] explicitly store past experiences and retrieve them via similarity measures. These methods, however, generally lack principled retrieval strategies and encounter difficulties in scaling to large memory sizes [ramani2019short]. Our framework addresses these issues by introducing density-based similarity priors and KL-regularised updates, which enable efficient and stable use of memory over time.
Case-based reasoning (CBR) [AamodtPlaza1994, Ashley1990] has also inspired memory-driven approaches to problem solving, emphasising the retrieval and adaptation of relevant past cases. In modern contexts, LLM agents increasingly employ explicit episodic memory systems to extend context and improve performance across diverse tasks [Zhong2024, Chhikara2025, fountas2024human]. Yet, the integration of memory retrieval with decision-making in LLMs still lacks a firm theoretical foundation, particularly regarding the conditions under which such systems converge to optimal behaviour. Our work addresses this gap directly.
@@
Recent work on Stable Hadamard Memory [hadamard2024] highlights the importance of stability in memory systems, using Hadamard products to calibrate and update stored information. While complementary in spirit, our method focuses instead on retrieval optimisation through information-theoretic regularisation. These two perspectives—stability of storage and optimality of retrieval—could be integrated in future research to build more comprehensive memory architectures.
@@
Finally, from a theoretical perspective, our framework extends established convergence results in reinforcement learning. The convergence proof for our soft policy iteration builds upon entropy-regularised MDPs [neu2017unified] but adapts them to the context of memory-based retrieval. Our two-time-scale analysis further draws on stochastic approximation theory [borkar2009stochastic], extending it to LLM-M-MDPs where value and memory are updated simultaneously. A central outcome is the formal equivalence we establish between retrieval, read, and write operations and the fundamental processes of policy improvement and policy evaluation. This connection makes retrieval-augmented LLMs provably learnable for the first time, opening the way for reinforcement learning theory and tools to be systematically applied in this emerging setting.
6 Conclusions
Our Parzen-KL framework makes several theoretical contributions. First, it introduces a principled notion of similarity, using Parzen windows to provide a statistically grounded approach to similarity-based retrieval. Second, it offers convergence guarantees, showing that both the soft and hard variants of the method converge under appropriate conditions. Third, it establishes reflection consistency, formally linking memory growth to progress toward optimality. Finally, it outlines practical algorithms, presenting clear implementation paths for both exact and approximate variants.

### v2

@@ 4.2 Two-Time-Scale Convergence with Online Memory Rewriting
    $\sum_{t=0}^{\infty}\eta_{t}=\infty,\quad\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty,\quad\sum_{t=0}^{\infty}\rho_{t}=\infty,\quad\sum_{t=0}^{\infty}\rho_{t}^{2}<\infty,\quad\lim_{t\to\infty}\frac{\rho_{t}}{\eta_{t}}=0.$
@@
    $\mathbb{E}[\xi^{(Q)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[\xi^{(\mu)}_{t}|\mathcal{F}_{t}]=0,\quad\mathbb{E}[(\xi^{(Q)}_{t})^{2}|\mathcal{F}_{t}]\leq K,\quad\mathbb{E}[(\xi^{(\mu)}_{t})^{2}|\mathcal{F}_{t}]\leq K.$
@@
    Memory attractor: The memory process $\{M_{t}\}$ has a compact attractor set $M_{\infty}\subset M$. This means memory size is bounded and the content distribution stabilises over time. In practice, this is achieved by: 1) Using fixed-size memory, 2) Having policies that converge to stationary distributions, 3) Using environments that produce stationary experience streams, or 4) Conservative memory update strategies.
@@
Under Assumptions 6 and 9, let $(Q_{t},\mu_{t},M_{t})$ be generated by the following sample-based recursions.
@@
Appendix C presents the required Lipschitz properties of the memory updates, and Appendix D provides the detailed proof of Theorem 10. The novelty lies in extending the classical ODE method to handle structured, state-dependent episodic memory updates and their coupling with fast policy dynamics under a two–time-scale regime. The key insight is that from the perspective of the fast variables $(Q_{t},\mu_{t})$, the slow variable $M_{t}$ appears quasi-static due to the timescale separation $\rho_{t}/\eta_{t}\to 0$. This detailed proof establishes several key properties of our Read–Write Learning iteration. First, it demonstrates robustness to memory updates, showing that the algorithm remains stable even as the memory evolves, provided that memory updates occur on a slower timescale than value and policy updates. Second, it guarantees tracking behaviour, ensuring that the value function and policy rapidly adjust to reflect changes in memory, remaining close to the optimal configuration for the current memory state. Third, the analysis confirms convergence to local optima, proving that for each fixed memory configuration (M), the algorithm converges to the corresponding optimal retrieval policy ($\mu^{*}_{M}$). Finally, the timescale separation result provides practical implementation guidance, indicating that memory should be updated less frequently than the value function and policy.
Algorithm 2 Read–Write Reflective Learning (Fast Policy Update, Slow Memory Write)
1:Input: Temperature $\alpha>0$, kernel bandwidth $h>0$, discount $\gamma\in(0,1)$
2:Stepsizes: $\{\eta_{t}\}$ for policy evaluation/improvement, $\{\rho_{t}\}$ for memory updates, with $\rho_{t}/\eta_{t}\to 0$, $\sum_{t=1}^{\infty}\eta_{t}=\infty$, $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$
3:Initialise: Memory $M_{0}$, Q-function $Q_{0}$, initial Parzen prior $\mu_{0,0}(\cdot\mid x)$ (incl. void case $c_{\varnothing}$)
4:for $t=0,1,2,\ldots$ do
5:  (1) State and Prior Construction
6:  Observe $s_{t}$ and form augmented state $x_{t}=(s_{t},M_{t})$.
7:  Construct the reference prior $\mu_{0,t}(\cdot\mid x_{t})$ from $M_{t}$ using (6) and (8).
8:  (2) Read: Policy Improvement (Fast Time Scale)
9:  Define the KL-greedy retrieval policy (Eq. 11) induced by $Q_{t}$:
$\mu_{t}(c\mid x_{t})=\frac{\mu_{0,t}(c\mid x_{t})\exp(Q_{t}(x_{t},c)/\alpha)}{\sum_{c^{\prime}}\mu_{0,t}(c^{\prime}\mid x_{t})\exp(Q_{t}(x_{t},c^{\prime})/\alpha)}.$
10:  Sample a retrieval action $c_{t}\sim\mu_{t}(\cdot\mid x_{t})$.
11:  (3) LLM Act
12:  Sample an environment action $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t})$ and execute it.
13:  (4) Environment Feedback
14:  Observe reward/feedback $r_{t}$ and next state $s_{t+1}$
15:  (5) a. Write (Slow Memory Update): update episodic memory,
$M_{t+1}\leftarrow(1-\rho_{t})M_{t}+\rho_{t}\,\mathrm{Write}(M_{t},s_{t},a_{t},r_{t},s_{t+1})\,$,
which denotes a generic slow update consistent with the assumed Lipschitz conditions.
16:  Next augmented state $x_{t+1}:=(s_{t+1},M_{t+1})$; Next prior $\mu_{0,t+1}(\cdot\mid x_{t+1})$ from $M_{t+1}$
17:   b. Write (Fast KL-regularised Evaluation): compute the target
$y_{t}:=r_{t}+\gamma\Bigg(\sum_{c^{\prime}\in\mathcal{C}(M_{t+1})}\mu_{t}(c^{\prime}\mid x_{t+1})\,Q_{t}(x_{t+1},c^{\prime})-\alpha\,\mathrm{KL}\!\big(\mu_{t}(\cdot\mid x_{t+1})\;\|\;\mu_{0,t+1}(\cdot\mid x_{t+1})\big)\Bigg),$
and update
$Q_{t+1}(x_{t},c_{t})\leftarrow(1-\eta_{t})Q_{t}(x_{t},c_{t})+\eta_{t}\,y_{t},$
which is stored in memory in the tabular case, or updated via function approximation.
18:end for
@@ 4.3 The Role of Reflection and Memory towards Optimal Policy
We are now ready to present the complete Read–Write Reflective Learning pseudocode in Algorithm 2, consistent with the workflow illustrated in Fig. 1.
To ensure practical feasibility, policy evaluation is implemented in an approximate form using feedback from the environment. Steps (2)–(4) and (5.b) constitute a fast policy iteration loop. The Read step performs a KL-regularised greedy update of the retrieval policy, while Step (5.b) approximately evaluates the resulting policy using feedback from the environment. The policy improvement is monotonic because the Read step solves a KL-regularised optimisation problem with respect to the current value function (Lemma 5). Given a fixed reference prior, the resulting update maximises the regularised expected return and therefore yields a policy that is no worse than the previous one in terms of the KL-regularised objective. Since the subsequent evaluation step provides an unbiased (or consistent) estimate of the policy value under environmental feedback, repeated application of Steps (2)–(4) and (5.b) leads to monotonic improvement of the retrieval policy on the fast time scale (Theorem 19). Step (5.a) implements the slow evolution of episodic memory. Under the two-time-scale assumptions, Theorem 10 establishes that the resulting algorithm converges to the optimal policy of the reflected MDP.
Crucially, episodic memory does not merely augment the state representation but directly parameterises the effective Bellman operator. Each memory update induces a new Reflected MDP and hence a new policy-evaluation problem. Although the $Q$-update takes the form of a standard temporal-difference step, the associated KL-evaluation operator is indexed by the current memory $M$ through the induced action set $\mathcal{C}(M)$, the memory-conditioned prior $\mu_{0}(\cdot\mid x,M)$, and the memory-dependent transition $x^{\prime}=(s^{\prime},\mathrm{Write}(M,\cdot))$.
Consequently, the read–write cycle realises a two-time-scale policy-iteration procedure: for fixed $M$, the effective MDP is stationary and the algorithm converges to the locally optimal policy induced by the current memory coverage, while memory growth drives successive improvements of the effective control problem. Since the Reflected MDP approximates the underlying environment MDP through LLM-mediated interaction (cf. (4)), a central question is whether a Read–Write Reflective Learning agent can asymptotically achieve optimal MDP performance as its episodic memory expands. We formalise this requirement via reflection consistency, which states that increasing memory coverage drives the composite policy—combining retrieval and LLM action selection—toward the true optimal policy.
@@
For an episodic memory $M\subset\mathfrak{M}$, define the coverage radius
$r_{M}:=\sup_{s\in\mathcal{S}}\min_{c\in M}d\bigl(s,s(c)\bigr),$
where $s(c)$ denotes the state component of memory case $c$. Equivalently, the statewise coverage radius is $r_{M}(s):=\min_{c\in M}d\bigl(s,s(c)\bigr)$.
Let the (deterministic) retrieval rule $c_{M}(s)$ pick any case achieving the minimum. The LLM-induced composite policy on the original state space is
$\pi_{M}(a\mid s)\;:=\;p_{\mathrm{LLM}}\big(a\mid s,c_{M}(s)\big).$
@@
The function $\varepsilon_{\mathrm{LLM}}(r)$ characterises the local decision-making capability of the LLM and can be interpreted as a measure of its local optimality gap. Specifically, it quantifies how closely the LLM-induced action distribution $p_{\mathrm{LLM}}(\cdot\mid s,c)$ approximates the optimal policy $\pi^{\star}(\cdot\mid s)$ when the retrieved memory case $c$ lies within a neighbourhood of radius $r$ around the true state $s$.
A small value of $\varepsilon_{\mathrm{LLM}}(r)$ indicates that the LLM is able to infer near-optimal actions using only coarse or approximate contextual information. In other words, even when the retrieved memory does not exactly match the current state, a strong LLM can leverage its internal semantic representations, world knowledge, and reasoning ability to interpolate correctly within a local region of the state space. Consequently, a good LLM is characterised by an error function $\varepsilon_{\mathrm{LLM}}(r)$ that remains small even for relatively large values of $r$, reflecting robustness to imprecise memory retrieval and limited state coverage.
Intuitively, $\varepsilon_{\mathrm{LLM}}(r)$ captures the effective “radius of competence” of the LLM: it measures the size of the neighbourhood within which the LLM can reliably recover near-optimal behaviour from local contextual cues. This property directly determines how densely episodic memory must cover the state space in order to achieve near-optimal performance. A smaller $\varepsilon_{\mathrm{LLM}}(r)$ for larger $r$ implies that fewer memory cases are required, leading to improved sample efficiency and faster convergence of the Read–Write Reflective Learning process.
We now examine the retrieval capability. For a given state $s$, define $\delta_{M}(s)$ as the probability that the retriever selects a case whose associated state lies outside radius $r_{M}(s)$, conditioned on the existence of at least one case within this radius, due to auxiliary retrieval errors (e.g., finite-bandwidth smoothing or finite-sample density estimation):
$\delta_{M}(s)\;:=\;\mathbb{P}\!\left(d\!\left(s,s(c)\right)>r_{M}(s)\;\middle|\;\exists\,c^{\prime}\in M:d\!\left(s,s(c^{\prime})\right)\leq r_{M}(s)\right).$
Lemma 12 (Retrieval Capability).
@@
$\delta_{M}\;:=\;\sup_{s\in\mathcal{S}}\delta_{M}(s).$
@@
$\Delta_{M}\;:=\;\sup_{s\in\mathcal{S}}\mathrm{TV}\!\big(\pi_{M}(\cdot\mid s),\pi^{\star}(\cdot\mid s)\big)\;\leq\;\varepsilon_{\mathrm{LLM}}(r_{M})+\delta_{M}.$
@@
Corollary 15 formalises the role of episodic memory in enabling asymptotically optimal behaviour without parameter updates. The result shows that the discrepancy between the memory-induced policy and the optimal policy of the underlying MDP decomposes into an LLM approximation error and a memory coverage error. As episodic memory grows and achieves sufficient coverage of the state space, in the sense that locally relevant memory cases enable the LLM to approximate optimal actions within neighbourhoods of each state, both error terms vanish. Consequently, the composite policy and its associated value function converge to the optimal solution. This establishes that iterative read–write reflection over expanding memory is sufficient to recover optimal performance in the limit.
@@ 5 Related Work
Our work on reflective memory dynamics and Read-Write Reflective Learning builds upon, yet also departs from, several established lines of research in reinforcement learning and large language models (LLMs). This section situates our contribution within the broader landscape of memory-based RL and clarifies the key distinctions from existing approaches.
The integration of memory mechanisms into reinforcement learning has long been explored to address the challenges of partial observability and sample complexity. Early approaches often relied on recurrent neural networks (RNNs), in particular Long Short-Term Memory (LSTM) networks [hochreiter1997long], to maintain hidden states that capture historical information. For instance, [hausknecht2015deep] extends DQN with an LSTM to summarise past observations for Q-learning under partial observability, providing an early neural-memory approach to POMDPs [kaelbling1998planning]. However, its memory is implicit and entangled in the recurrent hidden state, making long-horizon retention, interpretability, and revision of past information difficult.
@@
From another strand of memory-based RL, episodic control methods [ramani2019short] such as Model Free Episodic Control [blundell2016model] and Neural Episodic Control [pritzel2017neural] use memory primarily as a direct control mechanism, retrieving past transitions to approximate values or reuse actions via similarity matching. In this setting, memory effectively acts as a nonparametric policy, and generalisation is limited to local interpolation over stored experiences [ramani2019short].
In contrast, our Read–Write Reflective Learning uses episodic memory to enable reflection in large language models. Retrieved memory does not prescribe actions directly; instead, it provides local contextual grounding that allows the LLM to reason and generate actions within a semantically relevant neighbourhood. Memory thus serves as a substrate for reflective generalisation rather than as a policy itself. By combining density based similarity priors with KL regularised updates, our approach yields a principled, scalable retrieval mechanism that supports stable continual learning while exploiting the LLM’s inherent capacity for local semantic generalisation.
Recent work on Stable Hadamard Memory [hadamard2024] highlights the importance of stability in memory systems, using Hadamard products to calibrate and update stored information. While complementary in spirit, our method focuses instead on retrieval optimisation through information-theoretic regularisation. These two perspectives, stability of storage and optimality of retrieval, could be integrated in future research to build more comprehensive memory architectures.
Case-based reasoning (CBR) [AamodtPlaza1994, Ashley1990] has long inspired memory-driven problem solving through the retrieval and adaptation of past cases. Our prior work on CBR-based LLMs [guo2024ds, guo2025optimizing] provides empirical evidence, particularly in data science and coding tasks, that LLMs are highly effective at revising retrieved cases, enabling strong generalisation to new tasks and overcoming a critical bottleneck in traditional CBR systems. Within our framework, CBR-based LLMs can be viewed as a stateless special case, for which we provide a principled and unified treatment.
@@
Finally, from a theoretical perspective, our framework extends established convergence results in reinforcement learning. The convergence proof for our soft policy iteration builds upon entropy-regularised MDPs [neu2017unified] but adapts them to the context of memory-based retrieval. Our two-time-scale analysis further draws on stochastic approximation theory [borkar2009stochastic], extending it to SRDPs where value and memory are updated simultaneously. A central outcome is the formal equivalence we establish between read (retrieval) and write operations and the fundamental processes of policy improvement and policy evaluation. This connection makes retrieval-augmented LLMs provably learnable for the first time, opening the way for reinforcement learning theory and tools to be systematically applied in this emerging setting.
6 Conclusions and Future Work
In this paper, we have established a formal theoretical foundation for continual learning in LLM agents through episodic memory and reflection, without requiring parameter fine-tuning. The core contribution is the Stateful Reflective Decision Process (SRDP), which extends the classical MDP framework by modeling reflection as a dual-action process: a retrieval stage that reads from episodic memory (policy improvement) and an action generation stage conditioned on retrieved cases. By augmenting the state space to include memory, we prove this system is Markovian and can be analysed as a Reflected MDP. We then introduce Read-Write Reflective Learning, a practical algorithm that integrates Parzen-window-based retrieval with KL-regularised soft policy iteration. Key theoretical results include: (1) guaranteed convergence of the policy iteration algorithm under bounded rewards, (2) two-timescale convergence when memory evolves slowly relative to policy updates, and (3) asymptotic optimality: as episodic memory grows to cover the state space, the composite policy provably converges to the optimal policy of the underlying MDP. This framework unifies case-based reasoning [guo2024ds], retrieval-augmented generation, and reinforcement learning under a single mathematical formalism, providing the first rigorous convergence guarantees for memory-driven LLM agents.